If the lines whose equations are 3x+y-8=0 and -2x+by+9=0 are perpendicular, what is the value of b?

Answer by Alan3354(69443)

If the lines whose equations are 3x+y-8=0 and -2x+by+9=0 are perpendicular, what is the value of b?
-------------
Find the slope of 3x+y-8=0
y = -3x + 8
slope = -3
The slope of lines perpendicular is the negative inverse, m = +1/3
Put -2x+by+9=0 into slope-intercept form (that means solve for y)
by = 2x - 9
y = (2/b)x - 9/b
2/b is the slope, and it = 1/3
2/b = 1/3
b = 6